# Enhanced-FQL($\lambda$), an Efficient and Interpretable RL with novel Fuzzy Eligibility Traces and Segmented Experience Replay

### v1

@@ Abstract
This paper introduces a fuzzy reinforcement learning framework, Enhanced-FQL($\lambda$), that integrates novel Fuzzified Eligibility Traces (FET) and Segmented Experience Replay (SER) into fuzzy Q-learning with Fuzzified Bellman Equation (FBE) for continuous control tasks. The proposed approach employs an interpretable fuzzy rule base instead of complex neural architectures, while maintaining competitive performance through two key innovations: a fuzzified Bellman equation with eligibility traces for stable multi-step credit assignment, and a memory-efficient segment-based experience replay mechanism for enhanced sample efficiency. Theoretical analysis proves the proposed method convergence under standard assumptions. Extensive evaluations in continuous control domains demonstrate that Enhanced-FQL($\lambda$) achieves superior sample efficiency and reduced variance compared to n-step fuzzy TD and fuzzy SARSA($\lambda$) baselines, while maintaining substantially lower computational complexity than deep RL alternatives such as DDPG. The framework’s inherent interpretability, combined with its computational efficiency and theoretical convergence guarantees, makes it particularly suitable for safety-critical applications where transparency and resource constraints are essential.
@@ I INTRODUCTION
These challenges have spurred interest in alternative paradigms that maintain learning capability while offering greater suitability for embedded platforms. Fuzzy systems present a compelling alternative, combining inherent interpretability with built-in robustness to uncertainty through smooth membership functions and substantially lower parameter counts [15, 16]. However, existing fuzzy RL approaches themselves face limitations in scalability and learning efficiency that must be overcome. To address these limitations while preserving interpretability and efficiency, the Fuzzy Q-Learning framework introduced in [17] provides a baseline for integrating fuzzy inference with Bellman equilibrium to achieve robust and interpretable control. While effective in moderate-scale problems, this approach faces scalability limitations when applied to large state-action spaces, leading to slower convergence and lower sample efficiency [18, 19].
In this work, we enhance the Fuzzy Q-Learning structure with a novel fuzzified experience replay mechanism and segmented eligibility traces. This integration enables efficient multi-step credit assignment and better exploitation of past interactions, substantially improving stability and data efficiency. As a result, a more complete and scalable Enhanced Fuzzy Q($\lambda$)-learning approach is introduced, capable of handling complex tasks in continuous state–action space through efficient online learning.
The key contributions of this paper are summarized as: (I) As a novel Fuzzy–RL framework, we introduce the first integration of Fuzzified Bellman Equation (FBE) with a novel fuzzy eligibility traces, TD($\lambda$), and a segmented experience replay, enabling multi-step credit assignment and efficient value propagation across continuous state-action spaces. (II) By replacing deep neural networks with a structured fuzzy rule base, our method eliminates the need for complex architecture design and extensive hyperparameter tuning, drastically reducing the number of learnable parameters while maintaining approximation capability. (III) The smooth interpolation properties of the fuzzy inference system provide built-in robustness to input noise and sensor uncertainties, addressing a key limitation of neural network policies in safety-critical applications. (IV) A contraction mapping-based analysis of the proposed fuzzified Bellman operator is provided, ensuring convergence of the learning process and achieving the suboptimal policy.
@@ V SIMULATION RESULTS
For fuzzy approximation, Gaussian membership functions were employed to partition the state space into $N_{s}=\{3,3,7,5\}=315$ fuzzy sets and the action space into $N_{a}=5$ sets. The learning parameters are carefully tuned through experiments, learning rate $\alpha=0.005$, discount factor $\gamma=0.99$, trace decay $\lambda=0.8$, segment length $L=10$, batch size $B=32$, and exploration rate $\varepsilon$ decaying linearly from $0.2$ to $0.05$ over the first 500 episodes to balance exploration and exploitation.
To comprehensively evaluate the performance of the proposed Enhanced-FQL($\lambda$) method, a comparison study is designed.Comparison includes: a multi-step fuzzy Q-learning variant that utilizes $n$-step returns within the same fuzzy rule base, a fuzzy SARSA($\lambda$) approach serving as an on-policy counterpart that employs eligibility traces for policy evaluation, thereby highlighting the impact of the off-policy optimality operator utilized in our method, and DDPG that uses deep neural networks. This comparative implimentation enables a precise analysis of the individual contributions of each algorithmic component to the overall system performance.
Figure 1 illustrates the learning progression over $500$ training episodes, demonstrating the sample efficiency and superior performance of the proposed approach. Enhanced-FQL($\lambda$) achieved consistent convergence (to achieve better Return than -200) within approximately $129$ episodes, significantly outperforming n-step fuzzy Q-learning, fuzzy SARSA($\lambda$), and DDPG, baselines in terms of learning speed and final performance. All comparative evaluations were conducted under identical conditions, including fuzzy partitions, optimizer settings, exploration schedules, and similar random starting of the episode, to ensure a fair comparison.
Figure 1: Learning curves comparing Enhanced-FQL($\lambda$) with baseline methods on Cart&Pole benchmark. The proposed method demonstrates faster convergence and superior performance.
Table I provides a comprehensive quantitative comparison across multiple performance metrics. The proposed method achieved the highest average return. While DDPG delivered competitive performance, it required more computation time per update due to its neural network architecture.
@@
The comparison study validates the effectiveness of integrating fuzzy eligibility traces and segment-based experience replay into FQL–FBE. Three key observations emerge from our analysis:
First, the multi-step credit assignment through fuzzy eligibility traces ($\lambda=0.8$) significantly accelerated learning compared to one-step methods, reducing the sample requirements for convergence by approximately 35% compared to n-step FQL. This demonstrates the importance of temporal credit propagation in continuous control tasks.
Second, the segment-based experience replay mechanism effectively decorrelated training samples while maintaining temporal consistency, resulting in the lowest variance among all evaluated methods. This stability is particularly valuable for real-world applications where training data may be limited or expensive to acquire.
Third, while DDPG achieved respectable performance, Enhanced-FQL($\lambda$) provided a favorable trade-off between performance and computational efficiency. The fuzzy rule base required a few learnable parameters compared to thousands in neural network approaches, making it particularly suitable for resource-constrained embedded systems.
The inherent interpretability of the fuzzy rule base also allows for direct inspection of the learned control strategy, providing valuable insights into the decision–making process—a critical advantage in safety–sensitive applications where neural networks remain black boxes.
Figure 2 compares the speed of reaching a target performance (return = -200). Enhanced-FQL($\lambda$) achieves this with far fewer episodes, demonstrating its superior sample efficiency.
@@ VII CONCLUSION
This paper has introduced Enhanced-FQL($\lambda$), a novel fuzzy reinforcement learning framework that integrates new fuzzified eligibility traces, segmented experience replay, and a new fuzzy Q-learning architecture. The proposed method addresses key limitations of both deep reinforcement learning and conventional fuzzy RL approaches by offering a computationally efficient and interpretable solution for continuous control tasks. Theoretical analysis has established the convergence guarantee of Enhanced-FQL($\lambda$) under standard assumptions, while empirical results have demonstrated its superior sample efficiency and lower variance compared to n-step fuzzy TD and fuzzy SARSA($\lambda$) baselines. Notably, while achieving competitive performance with deep RL methods like DDPG, our approach maintains significantly lower computational requirements and provides inherent interpretability through its structured fuzzy rule base. The framework’s ability to handle continuous state-action spaces through fuzzy approximation, combined with efficient multi-step credit assignment and memory-aware experience replay, makes it particularly suitable for real-world applications where computational resources are constrained and safety-critical operation requires transparent decision-making.
Building upon these foundations, several promising research directions emerge from this study. First, the integration of prioritized experience replay (PER) could substantially enhance sample efficiency. Segment priority may be defined based on the average or maximum absolute TD error $\|\delta(\ell)\|$ across all activated fuzzy rules. Second, clustering-based experience replay could further optimize memory usage and learning speed and focus on more diverse or informative experiences, reducing redundancy and accelerating policy improvement. Third, extending the framework to multi-agent settings presents a valuable avenue. The interpretable nature of the fuzzy rule base and the sharing of tabular Q data could facilitate coordination and communication among agents. Fourth, exploring online rule adaptation—such as dynamically adjusting the number and placement of fuzzy sets ($N_{s}$, $N_{a}$, $c_{i}$, $\sigma_{i}$) based on incoming data—could enable automatic refinement of the state-action partition, balancing representation power and computational cost over time. Finally, incorporating model-based fuzzy reasoning may allow the agent to simulate and plan over fuzzy state transitions, combining the benefits of model-based reinforcement learning with the interpretability and robustness of fuzzy systems for more efficient long-horizon decision-making.

### v2

@@ Abstract
This paper introduces a fuzzy reinforcement learning framework, Enhanced-FQL($\lambda$), that integrates novel Fuzzified Eligibility Traces (FET) and Segmented Experience Replay (SER) into fuzzy Q-learning with the Fuzzified Bellman Equation (FBE) for continuous control. The proposed approach employs an interpretable fuzzy rule base instead of complex neural architectures, while maintaining competitive performance through two key innovations: a fuzzified Bellman equation with eligibility traces for stable multi-step credit assignment, and a memory-efficient segment-based experience replay mechanism for enhanced sample efficiency. Theoretical analysis proves the proposed method convergence under standard assumptions. On the Cart–Pole benchmark, Enhanced-FQL($\lambda$) improves sample efficiency and reduces variance relative to $n$-step fuzzy TD and fuzzy SARSA($\lambda$), while remaining competitive with the tested DDPG baseline. These results support the proposed framework as an interpretable and computationally compact alternative for moderate-scale continuous control problems.
@@ I INTRODUCTION
These challenges have motivated alternative paradigms that preserve learning capability while improving interpretability and implementation simplicity. Fuzzy systems offer a structured rule-based representation with smooth interpolation and relatively compact parameterization [15, 16]. However, existing fuzzy RL approaches still face limitations in scalability and learning efficiency. To address these limitations while preserving interpretability, the Fuzzy Q-Learning framework introduced in [17] provides a baseline for integrating fuzzy inference with Bellman equilibrium. While effective in moderate-scale problems, this approach faces scalability limitations as the state–action dimension and partition granularity increase, leading to slower convergence and higher computational cost [18, 19].
In this work, we enhance the Fuzzy Q-Learning structure with a fuzzified experience replay mechanism and segmented eligibility traces. This integration enables multi-step credit assignment and improved reuse of past interactions within a continuous state–action formulation. The resulting Enhanced Fuzzy Q($\lambda$)-learning approach is aimed at improving stability and sample efficiency while preserving interpretability in moderate-scale continuous-control settings.
The key contributions of this paper are summarized as: (I) We integrate the Fuzzified Bellman Equation (FBE) with fuzzy eligibility traces and segmented experience replay, enabling multi-step credit assignment within a continuous fuzzy state–action representation. (II) We formulate an interpretable rule-based alternative to neural function approximation for moderate-scale continuous-control problems. (III) We provide a contraction-based analysis of the proposed fuzzified Bellman operator, establishing convergence of the learning process and achieving the suboptimal policy. (IV) We validate the method on the Cart–Pole benchmark against $n$-step fuzzy Q-learning, fuzzy SARSA($\lambda$), and a DDPG baseline.
@@ V SIMULATION RESULTS
For fuzzy approximation, Gaussian membership functions were employed to partition the state space into $N_{s}=\{3,3,7,5\}=315$ fuzzy sets and the action space into $N_{a}=5$ sets. The learning parameters were tuned experimentally, with learning rate $\alpha=0.005$, discount factor $\gamma=0.99$, trace decay $\lambda=0.8$, segment length $L=10$, batch size $B=32$, and exploration rate $\varepsilon$ decaying linearly from $0.2$ to $0.05$ over the first 500 episodes. The segment length $L=10$ was selected as a practical compromise between preserving short-horizon temporal consistency for eligibility-trace reconstruction and maintaining sufficient decorrelation in replay; shorter segments truncate the effective trace more aggressively, whereas substantially longer segments reduce the replay benefit.
To evaluate the proposed Enhanced-FQL($\lambda$) method, a comparative study is conducted including: (i) a multi-step fuzzy Q-learning variant with $n$-step returns within the same fuzzy rule base, (ii) fuzzy SARSA($\lambda$) as an on-policy baseline with eligibility traces to assess the effect of the off-policy optimality operator, and (iii) DDPG using deep neural networks. For the DDPG baseline, the actor was implemented as a fully connected neural network with two hidden layers of 128 units each and ReLU activations, followed by a $\tanh$ output layer and a scaling layer mapping the action to the admissible force range. The critic was implemented as a state–action value network with separate state and action inputs, followed by two hidden layers of 256 units each with ReLU activations and a scalar output. Experience replay was used with a buffer capacity of $6\times 10^{5}$ transitions and a minibatch size of 256, with parameter updates starting after 5000 stored transitions. Target networks were updated using soft updates with coefficient $\tau=5\times 10^{-3}$. Exploration was induced by additive Gaussian action noise with initial standard deviation $1.5$, decayed multiplicatively by $0.997$, and actions were clipped to the admissible bounds. The discount factor was set to $\gamma=0.99$ and all results were averaged over 5 random seeds. This setup enables a clear analysis of the contribution of each algorithmic component to overall performance.
Figure 1 illustrates the learning progression over $500$ training episodes. Enhanced-FQL($\lambda$) reached the target return threshold in approximately $129$ episodes, converging faster than the fuzzy baselines and remaining competitive with the tested DDPG configuration on this benchmark. All methods were trained under the same environment, reward design, action bounds, episode horizon, random-start protocol, seed set, and evaluation procedure, while each baseline used its own algorithm-specific architecture and hyperparameterization.
Figure 1: Learning curves comparing Enhanced-FQL($\lambda$) with baseline methods on Cart&Pole benchmark. The proposed method demonstrates faster convergence and competitive performance.
Table I provides a quantitative comparison across multiple performance metrics. Enhanced-FQL($\lambda$) achieved the best average return among the evaluated methods and reached the target threshold earlier than the fuzzy baselines. The tested DDPG baseline delivered competitive final performance, but with higher per-update computation in the present implementation.
@@
The comparison study supports the effectiveness of integrating fuzzy eligibility traces and segment-based experience replay into FQL–FBE. Three observations are particularly relevant.
First, the multi-step credit assignment provided by fuzzy eligibility traces ($\lambda=0.8$) accelerated learning relative to the one-step fuzzy baseline, reducing the sample requirement for convergence by approximately 35% compared to $n$-step FQL.
Second, the segment-based experience replay mechanism effectively decorrelated training samples while maintaining temporal consistency, resulting in the lowest variance. This stability is particularly valuable where training data may be limited or expensive to acquire.
Third, the comparison with DDPG indicates that the proposed method can remain competitive on this benchmark while preserving an interpretable rule-based representation.
The fuzzy rule base also permits direct inspection of the learned control structure, which is a practical advantage when interpretability is important.
Figure 2 compares the speed of reaching a target performance threshold (return = -200). Enhanced-FQL($\lambda$) reaches this threshold in fewer episodes than the fuzzy baselines and earlier than the tested DDPG configuration on this benchmark.
@@ VII CONCLUSION
This paper introduced Enhanced-FQL($\lambda$), a fuzzy reinforcement learning framework that combines fuzzified eligibility traces, segmented experience replay, and a fuzzified Bellman backup within an interpretable rule-based representation. Theoretical analysis established convergence under standard assumptions to the fixed point induced by the selected fuzzy basis. On the Cart–Pole benchmark, the proposed method improved sample efficiency and reduced variance relative to the fuzzy baselines, while remaining competitive with the tested DDPG configuration. These results support Enhanced-FQL($\lambda$) as a promising interpretable alternative for moderate-scale continuous-control problems, while broader validation on more complex benchmarks remains an important direction for future work.